# Large Language Models for Cyber Security: A Systematic Literature Review

### v3

@@ 4.2. Trend Analysis
Illustrated in Figure 5, from 2020 to 2024, there have been significant shifts in the preference and utilization of LLM architectures across cybersecurity tasks. The selection of decoder-only, encoder-decoder, and encoder-only structures has influenced diverse research directions and solutions in the cybersecurity field. This examination delves into the trends regarding the adoption of these architectures over time, reflecting the evolving landscape of LLM applications for cybersecurity tasks.
@@
Timeline and Model Architecture distribution. In 2020 and 2021, the use of LLMs in cybersecurity was limited, with only 3 research papers exploring their potential. In 2020, encoder-decoder LLMs, known for their strong performance on sequence-to-sequence tasks, were the sole architecture used in a single paper. However, in 2021, the focus shifted to decoder-only LLMs, which excel at generating longer-form outputs and handling diverse queries due to their autoregressive generation capabilities and large context windows. This shift can be attributed to the research emphasis on LLM performance in natural language processing tasks and innovations in LLM architectures during this period (Kaplan et al., 2020; Brown et al., 2020).
@@
The year 2022 marked a significant turning point, with the number of papers employing LLMs for cybersecurity tasks surging to 11, surpassing the combined total from the previous two years. This year also saw increased diversity in the LLM architectures used. Encoder-only LLMs, valued for their representation learning and classification abilities, were utilized in 46% of the research (5 papers). Encoder-decoder LLMs, with their strong performance on well-defined tasks, were featured in 18% (2 papers), while decoder-only LLMs, leveraging their knowledge recall and few-shot learning capabilities, garnered 36% of the research interest (4 papers). This varied distribution highlights the active exploration of different architectures to address the diverse needs and challenges in cybersecurity.
@@
The years 2023 and 2024 witnessed a significant shift towards decoder-only LLMs, which emerged as the primary architecture for addressing cybersecurity challenges. This trend is closely tied to the powerful text comprehension, reasoning capabilities (Ouyang et al., 2022; Wei et al., 2023b), and open-ended generation demonstrated by chatbots like ChatGPT. These decoder-only models require minimal fine-tuning and can generate both syntactically correct and functionally relevant code snippets (Laskar et al., 2023a; Sadik et al., 2023). In 2023, decoder-only LLMs accounted for 68.9% of the total research, while encoder-decoder LLMs and encoder-only LLMs contributed 10.7% (14 papers) and 22.1% (27 papers), respectively. Remarkably, all studies conducted in 2024 utilized the decoder-only architecture, indicating a strong focus on exploring and leveraging the unique advantages of these models in cybersecurity research and applications.
@@
The dominance of decoder-only LLMs in cybersecurity research aligns with the broader trends in the LLM community. An analysis of the top 20 most liked and downloaded LLMs on Huggingface (hug, 2023), a popular open-source model community, reveals that while encoder-only models like BERT and its variants have the highest number of downloads, decoder-only models are gaining significant traction. Moreover, 16 out of the top 20 most liked LLMs are decoder-only models, indicating a strong preference and excitement for their potential to handle complex, open-ended tasks. The growing interest in decoder-only LLMs can be attributed to their strong generation, knowledge, and few-shot learning abilities, which make them well-suited for the diverse challenges in cybersecurity. However, the larger parameter size of these models compared to encoder-only models may limit their current adoption due to the scarcity of computational resources (Fedus et al., 2022).
Applying LLMs to cybersecurity. In our research, the use of LLMs can be categorized into agent-based processing and fine-tuning for specific tasks. Closed-source LLMs, represented by the GPT series, are the most popular in our studies. Researchers access LLMs online by calling APIs provided by LLM publishers and design task-specific prompts to guide LLMs to solve real-world problems with their training data (Jiang et al., 2020; Dong et al., 2023; Madaan et al., 2023), such as vulnerability repair and penetration testing (Deng et al., 2023a; Chen et al., 2023b; Wu et al., 2023a). Another approach involves locally fine-tuning open-source LLMs, by using datasets customized for specific functionalities, where researchers are able to achieve significant performance improvements (Silva et al., 2024; Yang et al., 2023c).
In summary, the transition of LLMs in cybersecurity, progressing from encoder-only architectures to decoder-only architectures, underscores the dynamic nature and flexibility of the field. This change has fundamentally altered the method for addressing cybersecurity tasks, signaling ongoing innovation within the discipline.
@@
RQ2 - Summary (1) We have gathered papers utilizing over 30 distinct LLMs for cybersecurity tasks. These LLMs have been categorized into three groups based on their underlying architecture or principles: encoder-only, encoder-decoder, and decoder-only LLMs. (2) We analyzed the trend in employing LLMs for cybersecurity tasks, revealing that decoder-only architectures are the most prevalent. Specifically, over 15 LLMs fall into the decoder-only category, and a total of 98 papers have investigated the utilization of decoder-only LLMs in cybersecurity tasks.
@@ 5.1. Fine-tuning LLMs for Security Tasks
Fine-tuning techniques are extensively utilized across various downstream tasks in NLP (Sun et al., 2020), encompassing the adjustment of LLM parameters to suit specific tasks. This process entails training the model on task-relevant datasets, with the extent of fine-tuning contingent upon task complexity and dataset size (Dodge et al., 2020; Qi et al., 2023b). Fine-tuning can mitigate the constraints posed by model size, enabling smaller models fine-tuned for specific tasks to outperform larger models lacking fine-tuning (Khare et al., 2023; Zheng et al., 2024b). We classify fine-tuning techniques employed in papers leveraging LLMs for security tasks into two categories: full fine-tuning and partial fine-tuning. Notably, many papers employ fine-tuning without explicitly specifying the technique. In such cases, if an open-source LLM is utilized, we presume full fine-tuning; if a closed-source LLM like GPT series models is utilized, we assume partial fine-tuning.
A total of 32 papers in this study applied fine-tuning techniques to address security tasks. Among them, the most popular approach is full fine-tuning, with 23 papers, accounting for 71.88% of the total. This may be because the pre-training tasks of LLMs are far from the content of the security tasks being applied, thus requiring updating all parameters of LLMs to achieve more competitive performance. partial fine-tuning is also highly regarded, with 28.12% of the papers choosing this approach to fine-tune LLMs. As shown in Table 7, full fine-tuning has a wide range of applications, including information and content security, network security, and software and system security. The most widespread domain among them is software and system security, totaling 16 papers, accounting for 65.57% of the total. A similar distribution is also observed in partial fine-tuning, with the most widely applied domain still being software and system security, totaling 7 papers, accounting for 77.78%. The applicability of fine-tuning techniques in these security tasks indicates that pre-training LLMs may not adequately address these security tasks, and updating model parameters on specific datasets is necessary to enhance effectiveness. The choice between full fine-tuning and partial fine-tuning depends on the balance between performance and efficiency considerations.
Table 7. Distribution of fine-tuning techniques adopted in papers and the numbers in parentheses represent the number of papers.
Fine-tuning technique Security task Reference
Full fine-tuning Bug detection (1) (Fang et al., 2023)
Access control (1) (Rando et al., 2023)
Steganography (1) (Wang et al., 2023b)
Reverse engineering and binary analysis (1) (Sun et al., 2023a)
Traffic and intrusion detection (1) (Ferrag et al., 2024)
Phishing and scam detection (2) (Ranade et al., 2021) (Jamal and Wimmer, 2023)
Harmful contents detection (2) (Hanley and Durumeric, 2023) (Mets et al., 2023)
System log analysis (2) (Karlsen et al., 2023) (Han et al., 2023)
Vulnerability repair (3) (Wu et al., 2023a) (Fu et al., 2022) (Zhang et al., 2023b)
Bug repair (4) (Yuan et al., 2022) (Paul et al., 2023) (Huang et al., 2023) (Wang et al., 2023a)
Vulnerability detection (5) (Chen et al., 2023b) (Ferrag et al., 2023) (Thapa et al., 2022) (Zhang et al., 2023g) (Khare et al., 2023)
Partial fine-tuning Traffic and intrusion detection (1) (Ali and Kostakos, 2023)
Harmful contens detection (1) (Hu et al., 2023b)
Program fuzzing (1) (Deng et al., 2023c)
Bug repair (2) (Jin et al., 2023) (Silva et al., 2024)
Bug detection (2) (Lee et al., 2023) (Yang et al., 2023c)
Vulnerability detection (2) (Chen et al., 2023c) (Khare et al., 2023)
Full fine-tuning. Full fine-tuning involves adjusting all parameters of the LLMs, including every layer of the model, to align with the specific requirements of the target task. This approach is favored when there exists a substantial disparity between the task and the pre-trained model or when the task necessitates the model to possess high adaptability and flexibility. Although full fine-tuning demands significant computational resources and time, it often yields superior performance (Lv et al., 2023). The success of full fine-tuning relies on having a dataset tailored to the task at hand. For instance, in bug fixing tasks, a dataset containing bug-patch pairs is essential to familiarize the LLMs with the intricacies of the target task (Huang et al., 2023; Wang et al., 2023a). LLM4Security encompasses a range of tasks, including bug repair (Huang et al., 2023; Wang et al., 2023a; Yuan et al., 2022; Paul et al., 2023), vulnerability detection and repair (Chen et al., 2023b; Zhang et al., 2023g; Fu et al., 2022), phishing, and harmful content detection (Mets et al., 2023; Jamal and Wimmer, 2023), among others, where full fine-tuning plays a crucial role in achieving optimal results.
Partial fine-tuning. Partial fine-tuning of LLMs is employed in some of the papers we collected, primarily to address security tasks while considering computational resource limitations and model copyright constraints. Partial fine-tuning involves updating only the top layers or a few layers of the model during the fine-tuning process, while keeping the lower-level parameters of the pre-trained model unchanged (Shen et al., 2021). The aim of this approach is to retain the general knowledge of the pre-trained model while adapting to the specific task by fine-tuning the top layers. This method is typically used when there is some similarity between the target task and the LLMs, or when the task dataset is small. In LLM4Security, the partial fine-tuning techniques applied can be categorized into API fine-tuning (OpenAI, 2023a; Hu et al., 2023b; Khare et al., 2023; Deng et al., 2023c; Jin et al., 2023; Ali and Kostakos, 2023), adapter-tuning (Houlsby et al., 2019; Lee et al., 2023; Yang et al., 2023c), prompt-tuning (Lester et al., 2021; Chen et al., 2023c), and Low-Rank Adaptation (LoRA) (Hu et al., 2021; Silva et al., 2024). These techniques ensure the effectiveness of LLMs in downstream security tasks while requiring smaller computational resource overhead.
@@ 5.2. Prompting LLMs for Security Tasks
Recent studies in natural language processing highlight the significance of prompt engineering (Liu et al., 2021) as an emerging fine-tuning approach aimed at bridging the gap between the output expectations of large language models during pretraining and downstream tasks. This strategy has demonstrated notable success across various NLP applications. Incorporating meticulously crafted prompts as features in prompt engineering has emerged as a fundamental technique for enriching interactions with large language models like ChatGPT, Bard, among others. These customized prompts serve a dual purpose: they direct the large language models towards generating specific outputs while also serving as an interface for tapping into the vast knowledge encapsulated within these models.
In prompt engineering, utilizing inserted prompts to provide task-specific knowledge is especially beneficial for security tasks with limited data features. This becomes crucial when conventional datasets (such as network threat reports, harmful content on social media, code vulnerability datasets, etc.) are restricted or do not offer the level of detail needed for particular security tasks. For example, in handling cyber threat analysis tasks (Siracusano et al., 2023), one can construct prompts by incorporating the current state of the network security posture. This prompts LLMs to learn directly from the flow features in a zero-shot learning manner (Xian et al., 2020), extracting structured network threat intelligence from unstructured data, providing standardized threat descriptions, and formalized categorization. In the context of program fuzzing tasks (Hu et al., 2023c), multiple individual test cases can be integrated into a prompt, assisting LLMs in learning the features of test cases and generating new ones through few-shot learning (Bendre et al., 2020), even with limited input. For tasks such as penetration testing (Deng et al., 2023a) and hardware vulnerability verification (Saha et al., 2023), which involve multiple steps and strict logical reasoning relationships between steps, one can utilize a chain of thought (COT) (Wei et al., 2023b) to guide the customization of prompts. This assists LLMs in process reasoning and guides them to autonomously complete tasks step by step.
@@ 5.3. External Augmentation
Features augmentation Incorporating task-relevant features implicitly present in the dataset into prompts. Adding bug descriptions, bug locations, code context or resampling for imbalanced traffic. (Jin et al., 2023) (Zhang et al., 2023a) (Xia et al., 2024) (Jamal and Wimmer, 2023)(Ali and Kostakos, 2023) (Zhang et al., 2023h) (Liu et al., 2023c)
External retrieval Retrieving task-relevant information available in external knowledge bases as input. An external structured corpus of network threat intelligence,a hybrid patch retriever for fix pattern mining. (Du and Yu, 2023) (Wang et al., 2023a)(Perrina et al., 2023)
External tools Analysis results from specialized tools serving as auxiliary inputs. Static code analysis tools, penetration testing tools. (Happe et al., 2023) (Alrashedy and Aljasser, 2024) (Armengol-Estapé et al., 2024)
Task-adaptive training Different training strategies from pre-training to enhance the model’s adaptability to the task. Contrastive learning, transfer learning, reinforcement learning, distillation. (Fang et al., 2023) (Lee et al., 2023)(Zhang et al., 2023b) (Pei et al., 2024)(Han et al., 2023) (Liang et al., 2024) (Tang et al., 2023)  (Cai et al., 2024)
Inter-model interaction Introducing multiple models (which can be LLMs or other models) to collaborate and interact. Multiple LLMs feedback collaboration, graph neural networks (Cai et al., 2024) (Yang et al., 2023a)(Tang et al., 2023)
Rebroadcasting Applicable to multi-step tasks, broad-casting the output results of each step iteratively as part of the prompt for the next step. Difficulty-based patch example replay, variables’ name propagation (Xu et al., 2023) (Yuan et al., 2022)
Post-process Customizing special processing strategies for LLMs’ outputs to better match task requirements. Post-processing based on Levenshtein distance to mitigate hallucinations, formal verification for generated code (Chen et al., 2023c) (Tihanyi et al., 2023)
@@
We summarized the external augmentation techniques combined with LLMs in previous studies, as shown in Table 8, with 7 different external augmentation techniques. The first augmentation technique we focus on is feature augmentation. The effectiveness of LLMs in handling downstream tasks heavily relies on the features included in the prompts. We have observed that many studies employing LLMs for security tasks extract contextual relationships or other implicit features from raw data and integrate them with the original data to customize prompts. These implicit features encompass descriptions of vulnerabilitiess (Zhang et al., 2023h), bug locations (Jin et al., 2023), threat flow graphs (Liu et al., 2023c), and more. Incorporating these implicit features alongside raw data leads to enhanced performance compared to constructing prompts solely from raw data. The next augmentation technique is external retrieval. External knowledge repositories can mitigate the hallucinations or errors arising from the lack of domain expertise in LLMs. LLMs can continually interact with external knowledge repositories during pipeline processing and retrieve knowledge relevant to security tasks to provide superior solutions (Perrina et al., 2023). Rule-based external tools can also serve as specialized external knowledge repositories. In addressing security tasks, LLMs can utilize results from external tools to rectify their outputs, thereby avoiding redundancy and errors(Happe et al., 2023; Alrashedy and Aljasser, 2024). The fourth augmentation technique is task-adaptive training. Existing studies adopt various training strategies from pre-training to strengthen LLMs’ adaptability to complex security tasks, enabling them to generate more targeted outputs. For instance, contrastive learning techniques can be employed, where both bugs and patches are used as input to LLMs to automatically generate higher-quality program patches (Chen et al., 2020; Tang et al., 2023). Alternatively, reinforcement learning can guide LLMs to produce more effective web test cases and alleviate local optima issues (François-Lavet et al., 2018; Liang et al., 2024). The fifth augmentation technique, inter-model interaction, has garnered significant attention when a single LLM may struggle to handle complex and intricate tasks. Decomposing the pipeline process and introducing multiple LLMs for enhanced performance have been explored (Yang et al., 2023a). This approach leverages collaboration and interaction among models to harness the underlying knowledge base advantages of each LLM. When a single interaction is insufficient to support LLMs in tasks such as variable name recovery or generating complex program patches  (Xu et al., 2023; Yuan et al., 2022), it is necessary to construct prompts for LLMs multiple times continuously to iterate towards the desired output. In this process, broadcasting the output results of each step iteratively as part of the prompt for the next step helps reinforce the contextual relationship between each interaction, thereby reducing error rates. The final augmentation technique is post-processing, where LLMs’ outputs are validated or processed for certain security tasks requiring specific types of output (Tihanyi et al., 2023). This process helps mitigate issues such as hallucinations arising from the lack of domain knowledge in LLMs (Chen et al., 2023c).

### v4

@@ 4.2. Trend Analysis
Illustrated in Figure 5, from 2020 to 2025 (data available up to early 2025), there have been significant shifts in the preference and utilization of LLM architectures across cybersecurity tasks. The selection of decoder-only, encoder-decoder, and encoder-only structures has influenced diverse research directions and solutions in the cybersecurity field. This examination delves into the trends regarding the adoption of these architectures over time, reflecting the evolving landscape of LLM applications for cybersecurity tasks.
@@
In 2020 and 2021, the use of LLMs in cybersecurity was limited, with only 1 and 2 instances identified respectively. In 2020, encoder-decoder LLMs were the sole architecture used (1 instance). However, in 2021, the focus shifted exclusively to decoder-only LLMs (2 instances). This early adoption pattern may be attributed to the research emphasis on LLM performance in natural language processing tasks and innovations in LLM architectures during this period (Kaplan et al., 2020; Brown et al., 2020).
@@
The year 2022 marked a significant turning point, with the number of instances employing LLMs for cybersecurity tasks surging to 11, surpassing the combined total from the previous two years. This year also saw increased diversity in the LLM architectures used. Encoder-only LLMs, valued for their representation learning and classification abilities, were utilized in 5 instances (45%). Encoder-decoder LLMs, with their strong performance on well-defined tasks, were featured in 2 instances (18%), while decoder-only LLMs, leveraging their knowledge recall and few-shot learning capabilities, garnered 4 instances (36%). This varied distribution highlights the active exploration of different architectures to address the diverse needs and challenges in cybersecurity.
@@
The years 2023 and 2024 witnessed a dramatic surge in research and a clear shift towards decoder-only LLMs becoming the predominant architecture. This trend is closely tied to the powerful text comprehension, reasoning capabilities (Ouyang et al., 2022; Wei et al., 2023b; Jiang et al., 2020; Dong et al., 2023; Madaan et al., 2023), and open-ended generation demonstrated by chatbots like ChatGPT. These decoder-only models often require minimal fine-tuning and can generate both syntactically correct and functionally relevant code snippets (Laskar et al., 2023a; Sadik et al., 2023). In 2023 (total 86 instances), decoder-only models constituted approximately 69% (59 instances), while encoder-only usage peaked at 23% (20 instances) and encoder-decoder models represented 8% (7 instances). This dominance of decoder-only models continued strongly into 2024 (total 85 instances), accounting for roughly 73% (62 instances). Importantly, unlike the suggestion of complete replacement, both encoder-decoder (12 instances,  14%) and encoder-only (11 instances,  13%) architectures continued to be utilized in 2024 across various domains, indicating their ongoing relevance for specific tasks. The limited data for early 2025 (6 instances) exclusively features decoder-only models, suggesting their continued prominence.
@@
The dominance of decoder-only LLMs in cybersecurity research aligns with the broader trends in the LLM community. An analysis of the top 20 most liked and downloaded LLMs on Huggingface (hug, 2023), illustrated in Table 6(b)), a popular open-source model community, reveals a divergence: encoder-only models like BERT and its variants represent the clear majority of the most downloaded models (13 out of 20), whereas decoder-only models overwhelmingly dominate the most liked list (16 out of 20). This indicates a strong community preference and excitement for the potential of decoder-only models to handle complex, open-ended tasks. The growing interest in decoder-only LLMs can be attributed to their strong generation, knowledge, and few-shot learning abilities, which make them well-suited for the diverse challenges in cybersecurity. However, the larger parameter size of these models compared to encoder-only models may limit their current adoption due to the scarcity of computational resources (Fedus et al., 2022).
In summary, while the cybersecurity field leverages all three main LLM architectures, recent years (2023-2024) have shown a clear dominance of decoder-only models, likely driven by their generative strengths and few-shot capabilities popularized by models like ChatGPT. However, the continued application of encoder-only and encoder-decoder models underscores their specific advantages for certain tasks, indicating a diverse rather than monolithic architectural landscape in LLM4Security.
@@
RQ2 - Summary (1) We have gathered papers utilizing over 30 distinct LLMs for cybersecurity tasks. These LLMs have been categorized into three groups based on their underlying architecture or principles: encoder-only, encoder-decoder, and decoder-only LLMs. (2) We analyzed the trend in employing LLMs for cybersecurity tasks, revealing that decoder-only architectures are the most prevalent. Specifically, over 15 LLMs fall into the decoder-only category, and a total of 133 papers have investigated the utilization of decoder-only LLMs in cybersecurity tasks.
@@ 5.1. Fine-tuning LLMs for Security Tasks
Fine-tuning techniques are extensively utilized across various downstream tasks in NLP (Sun et al., 2020), representing a crucial step in adapting pre-trained LLMs, whose initial training might not sufficiently cover the nuances of specialized domains like cybersecurity. This process involves further training the model on task-relevant datasets to adjust its parameters, aligning its capabilities more closely with specific security objectives (Dodge et al., 2020; Qi et al., 2023b). The extent and nature of fine-tuning often depend upon task complexity and dataset size (Dodge et al., 2020; Qi et al., 2023b). Importantly, fine-tuning can mitigate the constraints posed by model size, enabling smaller models fine-tuned for specific tasks to potentially outperform larger, general-purpose models lacking task-specific adaptation (Khare et al., 2023; Zheng et al., 2024b). In our surveyed literature (185 papers), fine-tuning was employed in 57 studies ( 31%). Its application varies significantly across domains: it was most prevalent in information and content security (10 out of 20 papers, 50%) and software and system security (43 out of 119 papers,  36%), with moderate use in network security (4 out of 26 papers,  15%). Notably, no papers applying fine-tuning were identified in the hardware security or blockchain security domains within our selection.
The primary motivation for fine-tuning in LLM4Security is to imbue general-purpose LLMs with domain-specific knowledge and capabilities, tailoring them for specific security tasks. For example, in software vulnerability management, models are fine-tuned to perform classification or generation. RealVul fine-tunes models like CodeBERT and CodeT5 on augmented datasets of PHP code snippets to improve vulnerability classification (Cao et al., 2024). VulLLM employs multi-task instruction fine-tuning on models including CodeBERT and CodeT5, training them simultaneously for vulnerability detection, localization, and explanation to enhance generalization (Du et al., 2024a). For vulnerability repair, fine-tuning on datasets containing vulnerable code and patches allows models like T5 (Hossain et al., 2024) or other code LLMs (Huang et al., 2023; Wang et al., 2023b; Wu et al., 2023a) to learn repair patterns. VulAdvisor further adapts this by fine-tuning CodeT5 with context-aware attention to generate natural language repair suggestions instead of direct code patches (Zhang et al., 2024d). Similarly, in information security, fine-tuning is applied to tasks like harmful content detection; EXPLAINHM fine-tunes a Flan-T5 model to classify harmful memes based on debates generated by a larger multimodal LLM (Lin et al., 2024). Fine-tuning is also used for adapting models to specific data types like network traffic (Ferrag et al., 2024) or system logs (Karlsen et al., 2023) for anomaly detection. This adaptation process typically involves preparing a high-quality dataset representative of the target task and using established training frameworks (Zheng et al., 2024b) to update the model’s parameters.
Different strategies exist regarding the scope of parameter updates during fine-tuning. Full fine-tuning adjusts all parameters of the LLM, offering the potential for high adaptability and performance, particularly when the target task differs significantly from the pre-training objectives (Lv et al., 2023). This approach is often employed when feasible, for instance, in tasks like bug repair (Huang et al., 2023; Wang et al., 2023b; Yuan et al., 2022; Paul et al., 2023) or vulnerability repair (Wu et al., 2023a; Fu et al., 2022; Zhang et al., 2023a) using dedicated datasets. However, full fine-tuning demands significant computational resources. Alternatively, Parameter-Efficient Fine-Tuning (PEFT) methods update only a small subset of parameters or add trainable adapters, significantly reducing computational and memory costs (Shen et al., 2021). PEFT techniques observed in LLM4Security include adapter-tuning (Houlsby et al., 2019; Lee et al., 2023; Yang et al., 2023b), prompt-tuning (Lester et al., 2021; Chen et al., 2023c), and Low-Rank Adaptation (LoRA) (Hu et al., 2021; Silva et al., 2024), alongside API-based fine-tuning for closed-source models (OpenAI, 2023a; Hu et al., 2024b; Khare et al., 2023; Deng et al., 2024c; Jin et al., 2023; Ali and Kostakos, 2023). These PEFT methods make fine-tuning more accessible, especially for very large models or in resource-constrained environments, while still achieving effective task adaptation.
Beyond standard supervised fine-tuning, Reinforcement Learning (RL) based strategies are gaining attention for aligning LLMs more closely with specific security objectives or leveraging their outputs to guide automated processes (Ouyang et al., 2022; DeepSeek-AI et al., 2025; François-Lavet et al., 2018). These approaches often involve optimizing an agent’s policy using reward signals derived from task execution or human feedback, sometimes incorporating LLM-generated insights into the reward function (Tang et al., 2024). For instance, in web application security, RL can fine-tune an LLM to generate more effective web fuzzing inputs by rewarding payload sequences that successfully bypass defenses or trigger desired responses (Liang et al., 2024). Similarly, RL techniques can enhance log analysis by optimizing models based on the accuracy or utility of the identified anomalies (Han et al., 2023). In the domain of fuzzing complex systems like JavaScript engines, CovRL utilizes coverage-guided RL: it processes execution coverage using methods like TF-IDF to assign higher importance to rarely executed paths, then uses this weighted coverage to generate a reward signal, thereby fine-tuning an LLM mutator to produce mutations more likely to explore these less-covered, potentially bug-rich areas (Eom et al., 2024). For automated program repair, RePair leverages RL with process-based feedback; instead of just learning from fixed code examples, the LLM’s policy for generating patches is directly rewarded based on the outcome of executing the generated code against test suites, thus encouraging the model to learn repair strategies that lead to functional correctness (Zhao et al., 2024). In a different application connecting LLMs and RL, CrashTranslator utilizes LLMs first to analyze stack traces and compute semantic similarities with UI elements; this information then shapes the reward function for an RL agent that learns an optimized UI exploration policy to effectively reproduce mobile application crashes (Huang et al., 2024).
Despite its benefits, fine-tuning LLMs for security tasks faces persistent challenges. Creating high-quality, large-scale, and representative labeled datasets remains a primary bottleneck; cybersecurity data is often scarce, imbalanced, noisy, or may contain inherent biases reflecting only known threats or specific environments, which can limit the model’s real-world applicability (Dodge et al., 2020). Furthermore, while generally less computationally intensive than pre-training, fine-tuning still demands considerable resources, particularly for full parameter updates on large models (Han et al., 2024). Performance limitations also exist. There is a significant risk of overfitting to the specific patterns present in the fine-tuning dataset, which can impair the model’s ability to generalize to novel or unseen vulnerability types, attack vectors, or code constructs encountered in practice (Montesinos López et al., 2022). Relatedly, catastrophic forgetting can occur, where the model loses some of its general language or reasoning capabilities acquired during pre-training (Luo et al., 2023). A critical concern, especially relevant to security, is that fine-tuning can inadvertently compromise safety alignments embedded during pre-training, potentially making models more susceptible to generating insecure code or harmful content if not carefully managed (Du et al., 2024b; Qi et al., 2023b).
@@ 5.2. Prompting LLMs for Security Tasks
Recent studies in natural language processing highlight the significance of prompt engineering (Liu et al., 2021) as an emerging fine-tuning approach aimed at bridging the gap between the output expectations of LLMs during pretraining and downstream tasks. This strategy has demonstrated notable success across various NLP applications. Incorporating meticulously crafted prompts as features in prompt engineering has emerged as a fundamental technique for enriching interactions with LLMs like ChatGPT, Bard, among others. These customized prompts serve a dual purpose: they direct the LLMs towards generating specific outputs while also serving as an interface for tapping into the vast knowledge encapsulated within these models.
In prompt engineering, utilizing inserted prompts to provide task-specific knowledge is especially beneficial for security tasks with limited data features. This becomes crucial when conventional datasets (such as network threat reports, harmful content on social media, code vulnerability datasets, etc.) are restricted or do not offer the level of detail needed for particular security tasks. For example, in handling cyber threat analysis tasks (Siracusano et al., 2023), one can construct prompts by incorporating the current state of the network security posture. This prompts LLMs to learn directly from the flow features in a zero-shot learning manner (Xian et al., 2020), extracting structured network threat intelligence from unstructured data, providing standardized threat descriptions, and formalized categorization. In the context of program fuzzing tasks (Hu et al., 2023b), multiple individual test cases can be integrated into a prompt, assisting LLMs in learning the features of test cases and generating new ones through few-shot learning (Bendre et al., 2020), even with limited input. For tasks such as penetration testing (Deng et al., 2024b) and hardware vulnerability verification (Saha et al., 2024), which involve multiple steps and strict logical reasoning relationships between steps, one can utilize a chain of thought (COT) (Wei et al., 2023b) to guide the customization of prompts. This assists LLMs in process reasoning and guides them to autonomously complete tasks step by step.
@@ 5.3. External Augmentation
Features augmentation Incorporating task-relevant features implicitly present in the dataset into prompts. Adding bug descriptions, bug locations, code context or resampling for imbalanced traffic. (Jin et al., 2023) (Zhang et al., 2024e) (Xia et al., 2024) (Jamal et al., 2024)(Ali and Kostakos, 2023) (Zhang et al., 2023f) (Liu et al., 2023c)
External retrieval Retrieving task-relevant information available in external knowledge bases as input. An external structured corpus of network threat intelligence,a hybrid patch retriever for fix pattern mining. (Du and Yu, 2023) (Wang et al., 2023b)(Perrina et al., 2023) (Wang et al., 2024d)(Wu et al., 2024b) (Zhang et al., 2024c)(Jiang et al., 2024) (Shan et al., 2024)(Yin et al., 2024b) (Liu et al., 2025)
External tools Analysis results from specialized tools serving as auxiliary inputs or validate the result with external tools. Static analysis tools, penetration testing tools, Rule-Based Reasoning Engine (Happe et al., 2023) (Alrashedy and Aljasser, 2024) (Armengol-Estapé et al., 2024) (Cao et al., 2024)(Nazzal et al., 2024) (Xu et al., 2025)(Chen and Jabbarvand, 2024) (Zhang et al., 2024c)(Zhang et al., 2024d) (Huang et al., 2024)(Zhang and Zhang, 2024) (Ahmad et al., 2024a)
Task-adaptive training Different training strategies from pre-training to enhance the model’s adaptability to the task. Contrastive learning, transfer learning, distillation. (Fang et al., 2023) (Lee et al., 2023)(Zhang et al., 2023a) (Pei et al., 2024)(Han et al., 2023) (Liang et al., 2024) (Tang et al., 2023)  (Cai et al., 2024a)
Inter-model interaction Introducing multiple models (which can be LLMs or other models) to collaborate and interact. Multiple LLMs collaboration, graph neural networks, gGAN (Cai et al., 2024a) (Yang et al., 2024a)(Tang et al., 2023) (Nazzal et al., 2024)(Huang et al., 2024) (Mündler et al., 2024)(Wang et al., 2024b)
Rebroadcasting Applicable to multi-step tasks, broad-casting the output results of each step iteratively as part of the prompt for the next step. Difficulty-based patch example replay, variables’ name propagation (Xu et al., 2023) (Yuan et al., 2022)(Xu et al., 2025)
Post-process Customizing special processing strategies for LLMs’ outputs to better match task requirements. Post-processing based on Levenshtein distance to mitigate hallucinations, formal verification for generated code (Chen et al., 2023c) (Tihanyi et al., 2023)(Lu et al., 2024) (Feng and Chen, 2024)(Wu et al., 2024c) (Wang et al., 2024b)
Workflow Integration Integrating LLMs as components within specific task workflows, incorporating various forms of interactive feedback. LLM-guided fuzzing strategy generation, Self-check reasoning, Coverage-guided feedback loop, Integrate LLM outputs into the human rating (Meng et al., 2024b) (Wang et al., 2024d)(Wang et al., 2024f) (Lyu et al., 2024)(Thomas et al., 2024) (Lin et al., 2024)
@@
We summarized the external augmentation techniques combined with LLMs in previous studies, as shown in Table 7, identifying 8 distinct categories. The first augmentation technique is feature augmentation. The effectiveness of LLMs in handling downstream tasks heavily relies on the features included in the prompts. We have observed that many studies employing LLMs for security tasks extract contextual relationships or other implicit features from raw data and integrate them with the original data to customize prompts. These implicit features encompass descriptions of vulnerabilities (Zhang et al., 2023f), bug locations (Jin et al., 2023), threat flow graphs (Liu et al., 2023c), and more. Incorporating these implicit features alongside raw data leads to enhanced performance compared to constructing prompts solely from raw data. The next augmentation technique is external retrieval. External knowledge repositories can mitigate the hallucinations or errors arising from the lack of domain expertise in LLMs. LLMs can continually interact with external knowledge repositories during pipeline processing and retrieve knowledge relevant to security tasks to provide superior solutions (Perrina et al., 2023). Next are external tools; rule-based external tools can serve as specialized external knowledge repositories, and in addressing security tasks, LLMs can utilize results from external tools (like static analyzers or formal methods) to rectify their outputs, thereby avoiding redundancy and errors(Happe et al., 2023; Alrashedy and Aljasser, 2024). The fourth augmentation technique is task-adaptive training. Existing studies adopt various training strategies (beyond standard fine-tuning discussed in Section 5.1), like contrastive learning or distillation, to strengthen LLMs’ adaptability to complex security tasks, enabling them to generate more targeted outputs (Chen et al., 2020; Tang et al., 2023; Fang et al., 2023). The fifth augmentation technique, inter-model interaction, has garnered significant attention when a single LLM may struggle to handle complex and intricate tasks. Decomposing the pipeline process and introducing multiple LLMs for enhanced performance have been explored (Yang et al., 2024a). This approach leverages collaboration and interaction among models to harness the underlying knowledge base advantages of each LLM. When a single interaction is insufficient to support LLMs in tasks such as variable name recovery or generating complex program patches  (Xu et al., 2023; Yuan et al., 2022), rebroadcasting is used: constructing prompts for LLMs multiple times continuously where the output results of each step are broadcast iteratively as part of the prompt for the next step. This helps reinforce the contextual relationship between each interaction, thereby reducing error rates. The next augmentation technique is post-processing, where LLMs’ outputs are validated or processed for certain security tasks requiring specific types of output (Tihanyi et al., 2023). This process helps mitigate issues such as hallucinations arising from the lack of domain knowledge in LLMs (Chen et al., 2023c). The last strategy is workflow integration, which involves integrating LLMs as components within specific task workflows, incorporating various forms of interactive feedback. This feedback, such as coverage metrics guiding fuzzing strategies (Wang et al., 2024f; Lyu et al., 2024), self-check reasoning steps (Wang et al., 2024f), or outputs being integrated into human rating systems (Thomas et al., 2024), allows the workflow’s context and objectives to steer the LLM’s behavior.